# The Monte Carlo Computational Summit – October 25 & 26, 2023 – Notre Dame, Indiana, USA

### v1

@@ Abstract
The Monte Carlo Computational Summit was held on the campus of the University of Notre Dame in South Bend, Indiana, USA on 25–26 October 2023. The goals of the summit were to discuss algorithmic and software alterations required for successfully porting respective code bases to exascale-class computing hardware, compare software engineering techniques used by various code teams, and consider the adoption of industry-standard benchmark problems to better facilitate code-to-code performance comparisons. A large portion of the meeting included candid discussions of direct experiences with approaches that have and have not worked. Participants reported that identifying and implementing suitable Monte Carlo algorithms for GPUs continues to be a sticking point. They also report significant difficulty porting existing algorithms between GPU APIs (specifically Nvidia CUDA to AMD ROCm). To better compare code-to-code performance, participants decided to design a C5G7-like benchmark problem with a defined figure of merit, with the expectation of adding more benchmarks in the future. Problem specifications and results will eventually be hosted in a public repository and will be open to submissions by all Monte Carlo transport codes capable of running the benchmark problem. The participants also identified the need to explore the intermediate and long-term future of the Monte Carlo neutron transport community and how best to modernize and contextualize Monte Carlo as a useful tool in modern industry. Overall the summit was considered to be a success by the organizers and participants, and the group shared a strong desire for future, potentially larger, Monte Carlo summits.
keywords:
Monte Carlo, neutronics, neutron transport, GPU, HPC, radiation transport, Exascale
articletype: CONFERENCE REPORT
@@ 1 Motivation
The last decade has seen significant efforts to develop Monte Carlo neutral particle transport codes to better utilize accelerated compute platforms (i.e., GPUs) (Pereira et al., 2013; Brantley et al., 2017; Hamilton and Evans, 2019), which represent the bulk of computational power in modern high-performance computers (HPC) and a significant amount in consumer-grade hardware  (Tramm, Yoshii, and Romano, 2023). Developing algorithms and methods that can take advantage of GPUs can provide significant speedups to users, allowing for more rapid prototyping and analysis of nuclear systems at ever greater fidelity.
The parallelism paradigm on GPUs is considerably different from that on traditional CPUs, meaning that algorithms that have been optimized for CPU-based performance have no guarantee of being equally efficient on GPUs. In fact, this has been the experience with some code bases: initially seeing slow down when first porting to GPUs (Pozulp et al., 2023). This situation is only further complicated by an API environment that is far more fractured and disparate than the codified one seen on CPUs with shared memory parallelism from OpenMP and distributed memory parallelism from MPI. Different vendors require the use of different proprietary toolchains that often do not translate well to other vendors’ hardware. Even support for seemingly basic operability on one vendor’s system can not be guaranteed on other systems. While there is support for open-source standards on multiple vendors’ hardware (e.g., OpenCL), the functionality of these libraries often lags critically behind vendors’ own APIs. As a result, effectively porting to GPUs can require making major alterations to both underlying algorithms and to software engineering structures of a given code. Different teams at different institutions have made different design choices for both basic algorithms and software engineering structures to achieve the same goal: computing ever more complex problems faster than ever.
As some codes have achieved the speedup promised by GPU-enabled stochastic transport others have had a more difficult time (Pozulp et al., 2023). Each software and algorithmic design choice has its advantages and concessions; while some initially promising results have provided (at best) lackluster performance improvements, others have been shown to be invaluable to Monte Carlo performance on GPUs. Finding a place where the Monte Carlo transport community could have frank discussions about what does and does not work was needed and provided the motivation for the Monte Carlo Computational Summit.
@@ 2 Introduction
[…]
List of participating institutions in alphabetical order and the number of people attending from each ($*{}^{*}$denotes CEMeNT university). Institution Number of Participants Argonne National Lab (ANL) 4 Dept. of Energy/NNSA 1 Lawrence Livermore National Lab (LLNL) 5 Los Alamos National Lab (LANL) 4 Massachusetts Institute of Technology (MIT) 1 Naval Nuclear Lab (NNL) 2 North Carolina State University ($*{}^{*}$ 2 Oak Ridge National Lab (ORNL) 5 Oregon State University ($*{}^{*}$ 3 Sandia National Lab (SNL) 1 Seattle University ($*{}^{*}$ 1 University of Illinois; Urban-Champagne (UIUC) 1 University of Notre Dame ($*{}^{*}$ 2 Total 31
@@ 3 Presentations
This section provides a brief overview and key highlights of the eight presentations given during the summit. Table 3 shows the title of each presentation along with the affiliated presenters, institutions, software, and ID listed in the presentation repository (McClarren and Morgan, 2023). Table 3 shows the names of the codes discussed, their availability, distribution license (if any), and primary repository URL if publicly available.
[…]
$*{}^{*}$resenter absent, ${}^{\dagger}$presentations not made available for inclusion in the repository. Title Presenter Affiliation Primary Code ID Advances in the Shift GPU solver Steven Hamilton ORNL Shift 2 Implicit Monte Carlo at LANL— Alex Long LANL Jayenne 3 Progress and Challenges MCATK GPU Prototype Code Tim Burke LANL MCATK ${}^{\dagger}$ LLNL Monte Carlo Transport Scott McKinley & LLNL Mercury 4 GPU Status Shawn Dawson High Energy Physics Monte Carlo Seth Johnson ORNL Celeritas 6 on GPU: Celeritas MC/DC+Harmonize: JIT Compilation Ilham Variansyah & CEMeNT MC/DC 5a & & Asynchronous GPU scheduling Braxton Cuneo (OSU/SU) 5b OpenMC at Exascale: Performance Portable John Tramm ANL OpenMC ${}^{\dagger}$ Monte Carlo Particle Transport on Intel, NVIDIA, and AMD GPUs Experience Adapting OpenMC for $*{}^{*}$ Gavin Ridley MIT OpenMC ${}^{\dagger}$ Monte Carlo Research at NNL Paul Burke NNL MC21 7
[…]
Software presented at the Monte Carlo Summit, their availability, licenses, and primary URLs. Software name Availability (license, if any) Primary URL Shift Distributed via RSICC rsicc.ornl.gov Jayenne Not publicly distributed MCATK Not publicly distributed Murcury Not publicly distributed OpenMC CPU Open source (specialized) github.com/openmc-dev/openmc OpenMC GPU Argonne Open source (MIT) github.com/exasmr/openmc OpenMC GPU MIT Open source (MIT) github.com/gridley/openmc/tree/cuda MC21 Not publicly distributed MC/DC Open source (BSD-3) github.com/CEMeNT-PSAAP/MCDC Harmonize Open source (BSD-3) github.com/CEMeNT-PSAAP/harmonize Celeritas Open source (Apache 2.0) github.com/celeritas-project/celeritas
@@ 3.1 Advances in the Shift GPU Solver
Shift is a continuous-energy CPU and GPU transport code, maintained at Oak Ridge National Laboratory (Pandya et al., 2016). Shift was originally written as a history-based CPU code with a rich set of features for modeling entire reactor systems. Many of these features have been ported to a separate GPU event-based code (Hamilton et al., 2021). Within the event-based scheme, the CPU essentially acts as an event dispatcher, calling events and passing the relevant data to the GPU. This means that the GPU and CPU back-ends share much of the same code, save for some data structure optimization and particle history re-implementation.
@@ 3.2 Implicit Monte Carlo at LANL—Progress and Challenges
Jayenne is a C++ implementation of the Implicit Monte Carlo (IMC) method described by Fleck and Cummings (1971) to solve thermal radiation transfer problems. Jayenne is called by the user using a Fortran API as part of a larger hydrodynamics code package (Thompson et al., 2021). The hydrodynamic “host” code passes in the relevant temperatures and opacities. Jayenne was first written at Los Alamos over 25 years ago, where it is still actively maintained and developed. This presentation covered the recent effort to port Jayenne from CPUs to GPUs, which required major refactoring.
@@ 3.3 MCATK GPU Prototype Code
Monte Carlo Application Toolkit (MCATK) is a Monte Carlo particle transport code from Los Alamos National Lab (Adams et al., 2015). The current GPU strategy involves moving the ray-tracing component of transport onto the GPU through a package called MonteRay (Sweezy, 2018). Ray tracing operations are batched then executed in unison on the GPU. In radiography applications (where there is a significant amount of streaming) they report an 80$\times$ performance increase. In simulations with few detectors they see a 1.5–2$\times$ increase in performance. MonteRay has single precision capabilities.
@@ 3.4 LLNL Monte Carlo Transport GPU Status
Mercury (Brantley et al., 2016) is a feature rich GPU- and CPU-capable particle transport code for neutrons, photons, and light element charged particles. Mercury has a sibling thermal radiative transport code, Imp (Brantley et al., 2019), which uses an implicit Monte Carlo method. Imp shares roughly 80% of its code with Mercury, so advances in hardware implementation provide benefits to both code bases. The LLNL Monte Carlo Transport Project has spent several years of development optimizing Mercury and Imp to meet the computing potential of new, GPU-centric architectures like Sierra and El Capitan. The earliest versions of the GPU code struggled to reach desired computational speedup values values (McKinley et al., 2019). Ongoing optimization efforts have resulted in improved GPU performance (Pozulp et al., 2023), although optimization efforts continue Currently, Mercury and Imp are performant on CPU and Nvidia architectures and are being extended to AMD architectures as El Capitan, an AMD-based machine at LLNL, is being delivered.
This architecture portability is achieved via “loop abstraction”, a process where macros are inserted around parallel loops. The source code in the loop is captured as a lambda expression. This lambda may be launched on the GPU with the standard chevron syntax, or on the host CPUs using OpenMP, or even run sequentially. The macro allows the compiler to generate code for all three scenarios. Where to execute is deferred to run time, and is controllable by the end user. This method also offers easy iteration and profiling during development, allowing programmers to simply modify a macro when implementing new ideas. This technique is similar to one used by RAJA (Beckingsale et al., 2019), a portability library developed at LLNL; while the loop abstraction macros may be ported to RAJA in the future, these methods were developed independently.
When porting their code to the GPU, the LLNL Monte Carlo Transport team decided to simply run their existing multiprocessor code on GPUs, use profiling to detect areas of low performance, and then optimize those areas. This method allowed for some developments that are rare among other transport codes, such as their use of managed memory within the GPU. Normally, managed memory allocation calls are considered too slow to be performant. However, Mercury and Imp use a library called Umpire (Beckingsale et al., 2020) to create a memory pool and amortize the cost of these calls, allowing managed memory to be performant.
@@ 3.5 High Energy Physics Monte Carlo on GPUs: Celeritas
Celeritas is a Monte Carlo transport code focused on finding a simulation that matches experimental detector output used in the upgraded High Luminosity Large Hadron Collider (HL-LHC) at CERN (Tognini et al., 2022). These simulations compute histories that represent analog physical realizations of a particle moving through the detector region of the HL-LHC, referred to as a “track”. Currently, these simulations use the Geant4 transport code (Agostinelli et al., 2003) to compute these realizations; however, the massive throughput of data from HL-LHC experiments requires accelerated computing on GPUs. Celeritas is designed to achieve this by re-structuring the data types and job kernels to run more efficiently on GPUs before passing them to Geant4. HL-LHC simulations are much more geometry intensive than neutron transport because of the high-energy charged particles being simulated, which requires the inclusion of magnetic fields, continuous slowing down, and multiple scattering. The simulations being run for the HL-LHC also require a higher degree of accuracy, at scales from millimeters to tens of meters; these problems combine to make these simulations a challenge to implement on GPU architectures.
Celeritas achieves this through a number of new implementations. Data structures are allocated in chunks on the CPU and passed to the GPU. This allows the GPU compiler to make better memory optimizations for the problem before running the main particle transport loop. This loop is written to handle large batches of tracks with an arbitrary number of kernels, using masking to hide unnecessary kernels on tracks where they are not utilized. This loop runs faster on a per-kernel basis, but sorting overhead reduces this speedup as the number of tracks increases. This also limits Celeritas to only tracking one set of particles at a time, an “event-granular parallelism” that can conflict with the track-granular parallelism expected by Geant4. The complex geometries used in these simulations cause Celeritas to run with far fewer particles on GPU due to implementation issues the VecGeom library (Wenzel, Sandro, Apostolakis, John, and Cosmo, Gabriele, 2020). The Oak Ridge Advanced Nested Geometry Engine (ORANGE) has also been considered to alleviate the geometry bottle neck on GPUs.
@@ 3.6 MC/DC$+$Harmonize: JIT Compilation & Asynchronous GPU scheduling
Monte Carlo Dynamic Code (MC/DC) is built and maintained by researchers associated with the Center for Exascale Monte Carlo Neutron Transport (Variansyah et al., 2023). MC/DC is fully open source (under a BSD-3 clause license) and written completely in Python with a compilation structure to support highly performant execution on CPU- and GPU-based architectures. It is written specifically to enable rapid deployment and testing of novel time-dependent transport algorithms and variance reduction techniques on heterogeneous architectures by a diverse user base in an academic setting.
@@
MC/DC couples a fully transient neutron transport algorithm to a novel (for the field) development scheme. In particle transport the physics kernels are compute kernels. This is different from other simulation types that can often rely on exterior packages to offload most of the compute complexity and handle optimizations for heterogeneous architectures (e.g., PetSC (Balay et al., 2023), MFEM (Anderson et al., 2021)). Thus, development of particle transport Monte Carlo codes has traditionally involved compiled languages due to speed limitations and requirements of heterogeneous architectures and their APIs. This can often make development difficult for users outside of a specific lab or field, or those developing on completely different systems, even if theoretically the same architecture. MC/DC remedies this by using an acceleration and abstraction scheme based on the Numba compiler for Python that supports compilation to CPU and GPU hardware targets (Lam, Pitrou, and Seibert, 2015). At runtime these compute kernels are lowered into intermediate representations in the LLVM compiler library and compiled to a specific hardware target. These compute kernels are then bound and executed by a larger Python script. The MC/DC team studied other acceleration and abstraction techniques (e.g., PyKokkos (Al Awar et al., 2021), PyOpenCL, and PyCUDA (Klöckner et al., 2012)) but determined that Numba was the least intrusive and easiest to use, while still providing significant speed-up for Monte Carlo applications (Morgan, Palmer, and Niemeyer, 2022).
MC/DC has a full unit, regression, and verification test suite built in and continuously integrated using GitHub Actions. One verification test is a time-dependent Kobayashi dog-leg benchmark problem (Kobayashi, Sugimura, and Nagaya, 2001). It uses the seven-group C5G7 cross section data (Hou et al., 2017) and a dog-leg duct filled with water. The initial condition is an isotropic pulse of particles in the highest-energy group. A fissionable region is located in the duct. The thermal flux is tallied in three dimensional space and tiem. In a strong-scaling analysis, the MC/DC team found that, after a constant 43 second compilation time, the Numba compiler provides a 462$\times$ speedup over pure Python mode.
@@
Harmonize (Cuneo and Bailey, 2023) is an asynchronous GPU scheduler that can take divergent code bases and dynamically reschedule when kernels get executed such that they binned together. It effectively becomes a wrapper around compute kernels to execute them in a way that decreases thread divergence on streaming architectures, thus increasing performance. From the Monte Carlo transport perspective this can be thought of as on-the-fly event-based scheduling; however, Harmonize can be generally implemented by any GPU program.
Using these acceleration and abstraction techniques, MC/DC has been able to go from nothing to a fully fledged transient continuous energy transport code with completely novel methods written by relatively novice developers in about two years. MC/DC developers have done successful explorations into quasi-Monte Carlo (Variansyah and McClarren, 2022b), hybrid iterative techniques for $k$-eigenvalue simulations (Pasmann et al., 2023b, a, 2024), transient population control techniques (Variansyah and McClarren, 2022a), automatic transient weight window production, hash-based random number generation, global uncertainty quantification (Clements et al., 2023), residual Monte Carlo methods, and machine learning techniques for dynamic node scheduling—among other active areas of transient methods exploration. As the benefit of some of these novel algorithms is proven in MC/DC, developers look to implement in other software packages, further pushing transient radiation transport methods into the future.
@@
OpenMC is a Monte Carlo full-featured neutron and photon transport code (Romano and Forget, 2013). For years, OpenMP threading has served as OpenMC’s shared-memory parallelism model, with MPI available for distributed memory parallelism. Tramm et al. (2022) recently refactored a branch of OpenMC to utilize OpenMP5 target offloading technology. This interface allows for shared-memory multiprocessing on multiple types of CPU and GPU platforms; as such, it is considered a “portable” interface that does not require modification to run on its supported architectures. This portability was the main motivation for the development of OpenMC, as there are many different architectures in use across the U.S. national laboratories, particularly in the realm of GPU programming. Device offloading was a new feature in OpenMP when development for OpenMC began; this was noted as a considerable risk for the development team, as offloading is used extensively in Monte Carlo codes. It was mentioned in discussion that this risk led to disappointing performance levels of several other OpenMP based (non-Monte Carlo) codes, and that OpenMC was fortunate enough to have a direct line of contact to OpenMP developers to help debug issues that arose during development, which aided in overcoming this risk.

### v2

@@ Abstract
The Monte Carlo Computational Summit was held on the campus of the University of Notre Dame in South Bend, Indiana, USA on 25–26 October 2023. The goals of the summit were to discuss algorithmic and software alterations required for successfully porting respective code bases to exascale-class computing hardware, compare software engineering techniques used by various code teams, and consider the adoption of industry-standard benchmark problems to better facilitate code-to-code performance comparisons. Participants reported that identifying and implementing suitable Monte Carlo algorithms for GPUs continues to be a sticking point. They also report significant difficulty porting existing algorithms between GPU APIs (specifically Nvidia CUDA to AMD ROCm). To better compare code-to-code performance, participants decided to design a C5G7-like benchmark problem with a defined figure of merit, with the expectation of adding more benchmarks in the future. The participants also identified the need to explore the intermediate and long-term future of the Monte Carlo neutron transport community and how best to modernize and contextualize Monte Carlo as a useful tool in modern industry. Overall the summit was considered to be a success by the organizers and participants, and the group shared a strong desire for future, potentially larger, Monte Carlo summits.
@@ 1 Motivation
The last decade has seen significant efforts to develop Monte Carlo neutral particle transport codes to better utilize accelerated compute platforms (i.e., GPUs) [1, 2, 3], which represent the bulk of computational power in modern high-performance computers (HPC) and a significant amount in consumer-grade hardware  [4]. Developing algorithms and methods that can take advantage of GPUs can provide significant speedups to users, allowing for more rapid prototyping and analysis of nuclear systems at ever greater fidelity.
The parallelism paradigm on GPUs is considerably different from that on traditional CPUs, meaning that algorithms that have been optimized for CPU-based performance have no guarantee of being equally efficient on GPUs. In fact, this has been the experience with some code bases: initially seeing slow down when first porting to GPUs [5]. This situation is only further complicated by an API environment that is far more fractured and disparate than the codified one seen on CPUs with shared memory parallelism from OpenMP and distributed memory parallelism from MPI. Different vendors require the use of different proprietary toolchains that often do not translate well to other vendors’ hardware. Even support for seemingly basic operability on one vendor’s system can not be guaranteed on other systems. While there is support for open-source standards on multiple vendors’ hardware (e.g., OpenCL), the functionality of these libraries often lags critically behind vendors’ own APIs. As a result, effectively porting to GPUs can require making major alterations to both underlying algorithms and to software engineering structures of a given code. Different teams at different institutions have made different design choices for both basic algorithms and software engineering structures to achieve the same goal: computing ever more complex problems faster than ever.
As some codes have achieved the speedup promised by GPU-enabled stochastic transport others have had a more difficult time [5]. Each software and algorithmic design choice has its advantages and concessions; while some initially promising results have provided (at best) lackluster performance improvements, others have been shown to be invaluable to Monte Carlo performance on GPUs. Finding a place where the Monte Carlo transport community could have frank discussions about what does and does not work was needed and provided the motivation for the Monte Carlo Computational Summit.
@@ 2 Introduction
Institution Number of Participants
Argonne National Lab (ANL) 4
Dept. of Energy/NNSA 1
Lawrence Livermore National Lab (LLNL) 5
Los Alamos National Lab (LANL) 4
Massachusetts Institute of Technology (MIT) 1
Naval Nuclear Lab (NNL) 2
North Carolina State University (NCSU) 2
Oak Ridge National Lab (ORNL) 5
Oregon State University (OSU) 3
Sandia National Lab (SNL) 1
Seattle University (SU) 1
University of Illinois; Urban-Champagne (UIUC) 1
University of Notre Dame (ND) 2
Total 31
Table 1: List of participating institutions in alphabetical order and the number of people attending from each (denotes CEMeNT university).
@@ 3 Presentations
This section provides a brief overview and key highlights of the eight presentations given during the summit. Table 2 shows the title of each presentation along with the affiliated presenters, institutions, software, and ID listed in the presentation repository [6]. Table 3 shows the names of the codes discussed, their availability, distribution license (if any), and primary repository URL if publicly available.
Title Presenter Affiliation Primary Code ID
Advances in the Shift GPU solver Steven Hamilton ORNL Shift 2
Implicit Monte Carlo at LANL— Alex Long LANL Jayenne 3
Progress and Challenges
MCATK GPU Prototype Code Tim Burke LANL MCATK N/A
LLNL Monte Carlo Transport Scott McKinley & LLNL Mercury 4
GPU Status Shawn Dawson
High Energy Physics Monte Carlo Seth Johnson ORNL Celeritas 6
on GPU: Celeritas
MC/DC+Harmonize: JIT Compilation Ilham Variansyah & CEMeNT MC/DC 5a &
& Asynchronous GPU scheduling Braxton Cuneo (OSU/SU) 5b
OpenMC at Exascale: Performance Portable John Tramm ANL OpenMC N/A
Monte Carlo Particle Transport on
Intel, NVIDIA, and AMD GPUs
Experience Adapting OpenMC for CUDA Gavin Ridley MIT OpenMC N/A
Monte Carlo Research at NNL Paul Burke NNL MC21 7
Table 2: Presenter absent, presentations not made available for inclusion in the repository.
Software name Availability (license, if any) Primary URL
Shift Distributed via RSICC rsicc.ornl.gov
Jayenne Not publicly distributed
MCATK Not publicly distributed
Murcury Not publicly distributed
OpenMC CPU Open source (specialized) github.com/openmc-dev/openmc
OpenMC GPU Argonne Open source (MIT) github.com/exasmr/openmc
OpenMC GPU MIT Open source (MIT) github.com/gridley/openmc/tree/cuda
MC21 Not publicly distributed
MC/DC Open source (BSD-3) github.com/CEMeNT-PSAAP/MCDC
Harmonize Open source (BSD-3) github.com/CEMeNT-PSAAP/harmonize
Celeritas Open source (Apache 2.0) github.com/celeritas-project/celeritas
Table 3: Software presented at the Monte Carlo Summit, their availability, licenses, and primary URLs.
@@ 3.1 Advances in the Shift GPU Solver
Shift is a continuous-energy CPU and GPU transport code, maintained at Oak Ridge National Laboratory [7]. Shift was originally written as a history-based CPU code with a rich set of features for modeling entire reactor systems. Many of these features have been ported to a separate GPU event-based code [8]. Within the event-based scheme, the CPU essentially acts as an event dispatcher, calling events and passing the relevant data to the GPU. This means that the GPU and CPU back-ends share much of the same code, save for some data structure optimization and particle history re-implementation.
@@ 3.2 Implicit Monte Carlo at LANL—Progress and Challenges
Jayenne is a C++ implementation of the Implicit Monte Carlo (IMC) method described by [9] to solve thermal radiation transfer problems. Jayenne is called by the user using a Fortran API as part of a larger hydrodynamics code package [10]. The hydrodynamic “host” code passes in the relevant temperatures and opacities. Jayenne was first written at Los Alamos over 25 years ago, where it is still actively maintained and developed. This presentation covered the recent effort to port Jayenne from CPUs to GPUs, which required major refactoring.
@@ 3.3 MCATK GPU Prototype Code
Monte Carlo Application Toolkit (MCATK) is a Monte Carlo particle transport code from Los Alamos National Lab [11]. The current GPU strategy involves moving the ray-tracing component of transport onto the GPU through a package called MonteRay [12]. Ray tracing operations are batched then executed in unison on the GPU. In radiography applications (where there is a significant amount of streaming) they report an 80$\times$ performance increase. In simulations with few detectors they see a 1.5–2$\times$ increase in performance. MonteRay has single precision capabilities.
@@ 3.4 LLNL Monte Carlo Transport GPU Status
Mercury [13] is a feature rich GPU- and CPU-capable particle transport code for neutrons, photons, and light element charged particles. Mercury has a sibling thermal radiative transport code, Imp [14], which uses an implicit Monte Carlo method. Imp shares roughly 80% of its code with Mercury, so advances in hardware implementation provide benefits to both code bases. The LLNL Monte Carlo Transport Project has spent several years of development optimizing Mercury and Imp to meet the computing potential of new, GPU-centric architectures like Sierra and El Capitan. The earliest versions of the GPU code struggled to reach desired computational speedup values values [15]. Ongoing optimization efforts have resulted in improved GPU performance [5], although optimization efforts continue Currently, Mercury and Imp are performant on CPU and Nvidia architectures and are being extended to AMD architectures as El Capitan, an AMD-based machine at LLNL, is being delivered.
This architecture portability is achieved via “loop abstraction”, a process where macros are inserted around parallel loops. The source code in the loop is captured as a lambda expression. This lambda may be launched on the GPU with the standard chevron syntax, or on the host CPUs using OpenMP, or even run sequentially. The macro allows the compiler to generate code for all three scenarios. Where to execute is deferred to run time, and is controllable by the end user. This method also offers easy iteration and profiling during development, allowing programmers to simply modify a macro when implementing new ideas. This technique is similar to one used by RAJA [16], a portability library developed at LLNL; while the loop abstraction macros may be ported to RAJA in the future, these methods were developed independently.
When porting their code to the GPU, the LLNL Monte Carlo Transport team decided to simply run their existing multiprocessor code on GPUs, use profiling to detect areas of low performance, and then optimize those areas. This method allowed for some developments that are rare among other transport codes, such as their use of managed memory within the GPU. Normally, managed memory allocation calls are considered too slow to be performant. However, Mercury and Imp use a library called Umpire [17] to create a memory pool and amortize the cost of these calls, allowing managed memory to be performant.
@@ 3.5 High Energy Physics Monte Carlo on GPUs: Celeritas
Celeritas is a Monte Carlo transport code focused on finding a simulation that matches experimental detector output used in the upgraded High Luminosity Large Hadron Collider (HL-LHC) at CERN [18]. These simulations compute histories that represent analog physical realizations of a particle moving through the detector region of the HL-LHC, referred to as a “track”. Currently, these simulations use the Geant4 transport code [19] to compute these realizations; however, the massive throughput of data from HL-LHC experiments requires accelerated computing on GPUs. Celeritas is designed to achieve this by re-structuring the data types and job kernels to run more efficiently on GPUs before passing them to Geant4. HL-LHC simulations are much more geometry intensive than neutron transport because of the high-energy charged particles being simulated, which requires the inclusion of magnetic fields, continuous slowing down, and multiple scattering. The simulations being run for the HL-LHC also require a higher degree of accuracy, at scales from millimeters to tens of meters; these problems combine to make these simulations a challenge to implement on GPU architectures.
Celeritas achieves this through a number of new implementations. Data structures are allocated in chunks on the CPU and passed to the GPU. This allows the GPU compiler to make better memory optimizations for the problem before running the main particle transport loop. This loop is written to handle large batches of tracks with an arbitrary number of kernels, using masking to hide unnecessary kernels on tracks where they are not utilized. This loop runs faster on a per-kernel basis, but sorting overhead reduces this speedup as the number of tracks increases. This also limits Celeritas to only tracking one set of particles at a time, an “event-granular parallelism” that can conflict with the track-granular parallelism expected by Geant4. The complex geometries used in these simulations cause Celeritas to run with far fewer particles on GPU due to implementation issues the VecGeom library [20]. The Oak Ridge Advanced Nested Geometry Engine (ORANGE) has also been considered to alleviate the geometry bottle neck on GPUs.
@@ 3.6 MC/DC+Harmonize: JIT Compilation & Asynchronous GPU scheduling
Monte Carlo Dynamic Code (MC/DC) is built and maintained by researchers associated with the Center for Exascale Monte Carlo Neutron Transport [21]. MC/DC is fully open source (under a BSD-3 clause license) and written completely in Python with a compilation structure to support highly performant execution on CPU- and GPU-based architectures. It is written specifically to enable rapid deployment and testing of novel time-dependent transport algorithms and variance reduction techniques on heterogeneous architectures by a diverse user base in an academic setting.
@@
MC/DC couples a fully transient neutron transport algorithm to a novel (for the field) development scheme. In particle transport the physics kernels are compute kernels. This is different from other simulation types that can often rely on exterior packages to offload most of the compute complexity and handle optimizations for heterogeneous architectures (e.g., PetSC [22], MFEM [23]). Thus, development of particle transport Monte Carlo codes has traditionally involved compiled languages due to speed limitations and requirements of heterogeneous architectures and their APIs. This can often make development difficult for users outside of a specific lab or field, or those developing on completely different systems, even if theoretically the same architecture. MC/DC remedies this by using an acceleration and abstraction scheme based on the Numba compiler for Python that supports compilation to CPU and GPU hardware targets [24]. At runtime these compute kernels are lowered into intermediate representations in the LLVM compiler library and compiled to a specific hardware target. These compute kernels are then bound and executed by a larger Python script. The MC/DC team studied other acceleration and abstraction techniques (e.g., PyKokkos [25], PyOpenCL, and PyCUDA [26]) but determined that Numba was the least intrusive and easiest to use, while still providing significant speed-up for Monte Carlo applications [27].
MC/DC has a full unit, regression, and verification test suite built in and continuously integrated using GitHub Actions. One verification test is a time-dependent Kobayashi dog-leg benchmark problem [28]. It uses the seven-group C5G7 cross section data [29] and a dog-leg duct filled with water. The initial condition is an isotropic pulse of particles in the highest-energy group. A fissionable region is located in the duct. The thermal flux is tallied in three dimensional space and time. In a strong-scaling analysis, the MC/DC team found that, after a constant 43 second compilation time, the Numba compiler provides a 462$\times$ speedup over pure Python mode.
@@
Harmonize [30] is an asynchronous GPU scheduler that can take divergent code bases and dynamically reschedule when kernels get executed such that they binned together. It effectively becomes a wrapper around compute kernels to execute them in a way that decreases thread divergence on streaming architectures, thus increasing performance. From the Monte Carlo transport perspective this can be thought of as on-the-fly event-based scheduling; however, Harmonize can be generally implemented by any GPU program.
Using these acceleration and abstraction techniques, MC/DC has been able to go from nothing to a fully fledged transient continuous energy transport code with completely novel methods written by relatively novice developers in about two years. MC/DC developers have done successful explorations into quasi-Monte Carlo [31], hybrid iterative techniques for $k$-eigenvalue simulations [32, 33, 34], transient population control techniques [35], automatic transient weight window production, hash-based random number generation, global uncertainty quantification [36], residual Monte Carlo methods, and machine learning techniques for dynamic node scheduling—among other active areas of transient methods exploration. As the benefit of some of these novel algorithms is proven in MC/DC, developers look to implement in other software packages, further pushing transient radiation transport methods into the future.
@@
OpenMC is a Monte Carlo full-featured neutron and photon transport code [37]. For years, OpenMP threading has served as OpenMC’s shared-memory parallelism model, with MPI available for distributed memory parallelism. [38] recently refactored a branch of OpenMC to utilize OpenMP5 target offloading technology. This interface allows for shared-memory multiprocessing on multiple types of CPU and GPU platforms; as such, it is considered a “portable” interface that does not require modification to run on its supported architectures. This portability was the main motivation for the development of OpenMC, as there are many different architectures in use across the U.S. national laboratories, particularly in the realm of GPU programming. Device offloading was a new feature in OpenMP when development for OpenMC began; this was noted as a considerable risk for the development team, as offloading is used extensively in Monte Carlo codes. It was mentioned in discussion that this risk led to disappointing performance levels of several other OpenMP based (non-Monte Carlo) codes, and that OpenMC was fortunate enough to have a direct line of contact to OpenMP developers to help debug issues that arose during development, which aided in overcoming this risk.